Grade 12th passMechanics

A monkey B is holding on to the tail of monkey A which is climbing up a rope.the masses of monkey A and monkey B are 5 kg and 2 kg respectively. if A can co-relatea tensionof 30 N in his tail then what force should it apply on the rope in order to carry monkey B with it. (take g=10 m/s^2)

subhash
8 Years agoGrade 12th pass

2 Answers

swarup kumar majee
8 Years ago
T-30-5g-5a=0                                         ------------1
30-2g-2a=0                                            ------------2
From equation 2 we get a=5m/s2
So,from equation one we get T-30-50-25=0  (as g=10m/s2 and a=5m/s2)
Therefore,T=105N
So,force exerted on the rope by monkey A in order to carry monkey B is 105N
Rishi Sharma
5 Years ago
Dear Student,
Please find below the solution to your problem.

T-30-5g-5a=0 ------------1
30-2g-2a=0 ------------2
From equation 2 we get a=5m/s2
So,from equation one we get T-30-50-25=0 (as g=10m/s2and a=5m/s2)
Therefore,T=105N
So,force exerted on the rope by monkey Ain order to carry monkey B is 105N

Thanks and Regards